1. In a math class, there are 12 boys and 14 girls. Five of the boys and 8 of the girls wear glasses. A person is chosen at random to be stuffed into cannon and fired through a flaming loop into a pool or tigers and crocodiles. What is the probability that
(a) The person wears glasses?
(b) The person is a boy or the person wears glasses?
(c) The person is a boy and the person wears glasses?
 

Answer by fcabanski(1391)

Probability = desired/total. There are 26 total students in the class.


A. There are 13 people who wear glasses. 13/26 = 1/2 = 50%


B. Boy or wears glasses is 12 boys, and 8 more (girls) who wear glasses for a total of 20. 20/26 = 10/13 = 76.9%


C. There are 5 people who are boys and wear glasses. 5/26 = 19.2%.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)